OpenStudy (anonymous):

Help Me Please I have 2min To get dis done or its and zero and then it will bring my grade down please help 2. The function y = –0.024t2 + 0.54t models the height y in feet of your pet frog's jump t seconds after it jumps. How far did the frog jump? How high did it go? (1 point) The frog jumped about 22.5 ft far and about 11.25 ft high. The frog jumped about 3.04 ft far and about 22.5 ft high. The frog jumped about 11.25 ft far and about 3.04 ft high. The frog jumped about 22.5 ft far and about 3.04 ft high. 4. Simplify the expression. (3 – i) – (2 + 6i) (1 point) 1 – 7i –1 + 7i –6i 5 + 5i What is the number of real solutions? –11x2 = x + 11 (1 point) cannot be determined one solution two solutions no real solutions

OpenStudy (amistre64):

3 points?

OpenStudy (anonymous):

4) (3 – i) – (2 + 6i) =3 – i – (2 + 6i) =3 – i – 2 - 6i add and subtract =? is it = 1-7i ? yes or no?

OpenStudy (anonymous):

i got it yall thanks for helping

OpenStudy (anonymous):

–11x^2 = x + 11 0=11x^2 + x + 11 using quadratic formula it ha s sqrt (- number) therefore the solution are pair of imaginary or complex numbers

OpenStudy (anonymous):

or no real number solutions

OpenStudy (anonymous):

on 2) y = –0.024t2 + 0.54t =0 factor it to get t=0 ,t=0.54/0.024=? is it 22.5? also use derivative then set it to zero, to get the max or highest jump y = –0.024t2 + 0.54t dy/dt = 2(–0.024)t + 0.54 =0 what is t? is it 11.25 sec if t=11.25 sec sub it to the formula y = –0.024t2 + 0.54t y=3.04 ? yes or no?